# Benchmarking Quantum Computers: Towards a Standard Performance Evaluation Approach

### v3

@@ II.2 What types of benchmarks are there?
Benchmarks can be classified according to multiple criteria, ranging from the objective of what portion of the behavior space they want to assess to the way the workload is defined. All of these categories fall under one of the following three different strategies when trying to assess the performance Lilja2005o ; Kounev2020mh :
@@
  • Fixed-work workloads: This benchmarking strategy refers to the types of benchmarks that focus on measuring the amount of time required to perform a fixed amount of computational work, as seen in the case of the old classic Dhrystone and Whetstone benchmarks. These benchmarks compiled a list of programs to be executed, the former favouring floating point operations, while the latter focuses on integer performance analysis WhetstoneBook ; DhrystonePaper .
@@
  • Fixed-time workloads: This benchmarking strategy refers to the types of benchmarks that focus on measuring the amount of computation performed within a fixed period, as seen in the case of the SLALOM benchmark. This benchmark runs an algorithm to calculate the radiosity, with the performance metric being the accuracy of the answer computed within a fixed time SLALOMBenchmark .
@@
  • Variable-work and variable-time workloads: This benchmarking strategy refers to the types of benchmarks that allow both the amount of computation performed and the time to vary. A great example is the HINT benchmark that aimed to find rational upper and lower bounds for $\int_{0}^{1}\frac{1-x}{1+x}.\,dx$. The assessment was later performed by calculating the quality improvements per second (QUIPS), i.e. dividing the quality of the solution by the execution time GustafsonAndSnell_HINT .
@@
In general, the performance of a system depends on the performance of multiple components of the system that are utilized during a certain operation. The problem with using fixed-work benchmarks is that they introduce an intrinsic performance bottleneck when measuring the increment of improving a single component of the system. Benchmarking models have been developed changing the workload type precisely to address this concern, and further discussions on overcoming issues with the fixed-work via using fixed-time benchmarks can be found in Ref. Gustafson1992 ; Isoefficiency . However, benchmarks that allow for variability in both time and workload provide greater flexibility to evaluate the performance of different systems, determining which machine produces the highest-quality result in the shortest time possible.
@@
One approach to benchmarking classification is obtained by focusing on the level of implementation required from usersHowToBuildABenchmark ; Kounev2020mh :
@@
The specification-based benchmarks are sometimes called ”pen and paper” benchmarks, and they have proved successful when there are considerable differences in the architectures on which the benchmarks will be implemented ReflectionsOnCreationParallelBenchmarkSuite . Still, they come at the cost of ad-hoc and expensive implementations from their users, as well as challenges with verifiability. On the other hand, kit-based benchmarks reduce the cost and time to run them by offering execution-ready implementation, which favoured them historically. Indeed, the stone aged benchmarks for classical computing are good examples of this, such as the Whestone, Dhrystone or the Linpack benchmarks WhetstoneBook ; DhrystonePaper ; OverviewOfCommonBenchmarks .
Another approach to benchmarking classification is obtained by focusing on the workload Kounev2020mh ; Lilja2005o :
@@
II.3 What makes a good metric
@@
From Lilja and Kounev Lilja2005o ; Kounev2020mh , we know that these can be typically classified into three groups: Counting how often a certain event occurs (e.g. number of operations processed in a given observation interval), duration of an interval (e.g. time taken to complete a single operation), and size of a parameter (e.g. precision obtained to compute a value). Some examples of basis performance metrics are the response time (i.e., time taken by a system to react to a request providing a respective response), throughput (i.e., rate at which requests are processed by a system), and utilization (i.e., fraction of time in which a resource is used).
Similarly to the quality attributes defined for a benchmark, there are properties proposed in the literature for an ideal performance metric Kounev2020mh ; Lilja2005o ; HowToBuildABenchmark :
@@
We can also classify metrics depending on what they measure. Metrics that measure what was done, useful or not, have been called means-based metrics whereas end-based metrics measure what is actually accomplished. As an example, for a given program we could measure the floating point operations per second (FLOPS), or the total execution time. The former is a mean-based metric, while the latter is an end-based metric. Depending on the context, it could be argued that one is better than the other Lilja2005o .
@@
Other well-known examples of classical metrics are clock rate, or million instructions per second (MIPS). The clock rate is repeatable (given a fixed processor), practical, consistent, and independent metric. However, given its limited scope, the clock rate is both non-linear and unreliable as a performance metric (a processor with a faster clock rate does not necessarily imply better overall performance. The clock rate ignores many important performance-relevant aspects such as how much computation is actually performed in each clock cycle as well as the interaction with the memory and I/O subsystems). On the other hand, MIPS is neither reliable, linear, or consistent (for the latter, take the example of different Instruction Set Architectures, such as RISC vs CISC IsaWars ). Ultimately, we mentioned before FLOPS. Unfortunately, this metric is neither reliable, linear, consistent, nor independent (processor manufactures may use different rules for what is counted as floating-point operations, which makes the metric fail for consistency and independence).
@@
Finally, it is important to highlight two key points. First, the properties discussed in this work are applicable not only to full-stack benchmarks and metrics but also to individual components of the computing stack. Second, the seemingly straightforward task of measuring a property using a metric might be tough. The effect of the Simpsons’ paradox serves as an adage to the designers of benchmarks and metrics, calling out the possibility of reversing or disappearing the trend on the data when combining it leboudec2010performance .
II.4 Examples of classical benchmarks
Classical computing benchmarking has evolved to deal with multiple and distinct applications, from Machine Learning processing to High Performance Computing (HPC). In each of these fields, ad-hoc efforts are required to define the proper benchmarks. As an example, the original LINPACK benchmark based on subroutines for analyzing and solving systems of simultaneous linear algebraic equations and linear least squares problems has been used as the base for the HPLinpack highly parallel computing benchmark for HPC evaluation dongarra1979linpack ; Dongarra2003TheLB . Indeed, the Top 500 list top500 ranks the most powerful non-distributed computer systems using HPLinpack. Interestingly, there are still discussions about the suitability of this approach for HPC benchmarking CriticsTop500 . For instance, the HPLinpack benchmark overrates the available compute cycles for real-world applications. Consequently, alternatives such as the High Performance Conjugate Gradients (HPCG) benchmark are being explored as an alternative for evaluating supercomputers for the TOP 500 list (unfortunately, even the alternatives such as the HPCG are not immune to techniques that artificially inflate the problem size to rig the results) PerformanceModelingHPCG .
Other examples of controversial benchmarks used historically are the synthetic ones, such as the Whetstone WhetstoneBook . Even though it was proposed to measure the floating point performance of a CPU, it can be misleading when compared with real-life situations. Nowadays, computers can easily fit the main loop of Whetstone in the low-level caches of the CPU, keeping the math coprocessor pipeline filled and achieving its optimal speed. Moreover, they face the issue of most synthetic benchmarks, they are not guaranteed to be correlated with real application scenarios, and fail to meet true relevance.
The issues with Whetstone as well as LINPACK showed the need for updating and adapting benchmarks to keep up with the advances in technology. A possible way to deal with this is the approach taken by the benchmarks developed by Standard Performance Evaluation Corporation. These benchmarks are updated regularly. What is more, they have been developing custom-made benchmarks for differents types of applications. For instance, SPECviewperf 2020 allows users to measure graphics performance gwpg.spec.org while SPECaccel 2023 helps when assessing the performance of computationally intensive parallel applications running under the OpenACC and OpenMP target offloading APIs spec.org.accel2023 , or SPEC CPU 2017 that enables comparison of compute-intensive workloads resolution capacity on different computer systems (including optional metrics to measure energy consumption)spec.org.cpu2017 .
III Quantum benchmarking
@@
Quantum computers are devices that leverage the principles of quantum mechanics to address challenges beyond the computational capacity of classical computers. Several physical implementations for quantum computers have been explored, each employing different technologies as qubits, such as trapped ion QCion ; Zhang_2017 , superconducting circuits Arute2019 ; ZHU2022240 ; PhysRevLett.127.180501 , photonic QCPhoton ; PhysRevLett.127.180502 , and cold atoms Bernien2017ProbingMD ; Ebadi2020QuantumPO . Each of these architectures has its own set of advantages and disadvantages, and research is ongoing to determine which approach is most suitable for scaling up quantum computers to a practical level. Additionally, various quantum computing paradigms have emerged, including adiabatic quantum computing, which relies on adiabatic theorem to solve computational problems farhi2000 ; analog quantum computing, in which a controllable quantum device used to simulate other quantum systems Kendon_2010 ; digital quantum computing, which relies on the quantum circuit model for implementing general quantum algorithms doi:10.1126/science.273.5278.1073 ; and measurement-based quantum computing, where computation is performed through a sequence of measurements on a highly entangled state Raussendorf . These four paradigms offer different perspectives on the practical applications of quantum computing.
@@
The diversity of technological platforms and paradigms used in the development of quantum computers introduces significant complexity in comparing the performance among different devices. For this reason, the search for a common benchmark that allows comparison of quantum computers performance across different technological platforms and paradigms poses a challenge to the scientific community. Despite these challenges, efforts have been made to develop benchmarks enabling comparison among diverse quantum computers. Primarily, these efforts focus on three types of relevant comparisons: comparing the performance of different quantum processors within the same computational paradigm (e.g., comparison of digital quantum computing platform providers such as Forest, Qiskit, ProjectQ LaRose2019overviewcomparison , or comparisons between quantum processors Linke2017 ; miessen2024benchmarkingdigitalquantumsimulations ; Ilin_2024 ; wright2024noisyintermediatescalequantumsimulation ); comparing different computational paradigms (e.g., between processors for digital quantum computing and quantum annealing or digital-analog quantum computing Flannigan_2022 ); and finally, comparing quantum computers to classical computers (e.g. comparing digital processing against quantum annealing Jnger2021 , empirical comparisons relying upon simplified Machine Learning (ML) tasks havenstein2018comparisons ; Cugini2023 or even the claims on quantum supremacy Arute2019 ; kalai2022google ). It is worth noting that these three comparisons are crucial for advancing scalable and reliable quantum technologies, but each of them demands a thorough analysis. Furthermore, a similar analysis could be expanded to include algorithmic comparisons. Looking ahead, benchmarking should be extended to cover other potential new aspects of a quantum computer that we have not even envisioned yet (e.g. qRAM QuantumRam or other future components that we have not even thought about yet). During this work, we will restrict ourselves to the evaluation of the performance of quantum processors.
@@
Despite technological advances, we are still undergoing the development of the quantum computing foundations. As Blume-Kohout and Young stated osti_1594677 , we are still more in the early stages of the quantum computing technology world rather than close to any exponential growth, which translates into challenges as we do not have clearly defined applications for these devices. What is more, we lack knowledge of how each the performance of each component will propagate and how the different quantum technology architectures will evolve. The current state is closely related to classical devices in the pre-1940 stages with the introduction of devices that were able to mechanize sequential series of calculations and procedures in binary code via the Bell Labs Model I, which served as an example of successfully applying Boolean logic to the design of computers TROPP1980115 . Back then, computers were starting to be explored, in contrast to upcoming years such as 1965 when Moore published the famous paper that led to the establishment of the Moore’s law MooreLaw , and exponential growth took off. This contrasts with the ”benchmarketing” approach that we sometimes see in the community when promoting the results of quantum computing experiments.
@@
III.1 Nuances of benchmarking quantum computers
There is lack of a definite computational paradigm, standard, platform, or even architecture for quantum computers. This is an eventual consequence of our limited knowledge so far about the real scope of quantum computing devices and platforms. Note that, in classical computing, regardless of the type of architecture (i.e. Princeton/Von Neumann vs. Harvard), there is a similar unified approach that allows for performance evaluation using well-established metrics, workloads, and even benchmarks. Indeed, it can be argued that differences in the architectures have been historically deceptive mythHarvard , and that devices can be easily compared assuming a similar common ground. Interestingly, this was not always the case. As mentioned before, back in the times of the Bell Labs Model I, every computer device was custom and hard to compare against each other. Benchmarking was not so extended and generalised as a recommended best practice until the 70s-80s.
@@ III.1 Nuances of benchmarking quantum computers
    Noisy Intermediate-Scale Quantum (NISQ) Preskill2018 , characterized by middle-size quantum processors with tens to hundreds of noisy physical qubits, substantially limiting the size of the quantum circuits that can be reliably executed. In the current state of NISQ era, it is controversial whether quantum processors have already reached the necessary technological development to achieve quantum advantage.
@@
    Partially Quantum Error Corrected (PQEC), when quantum processors errors are below a certain threshold that allows the application of error correction codes, such as the surface code Fowler2012 . When an insufficient number of physical qubits are available, we could employ a smaller code and partially correct errors up to a certain distance-code and number of errors per per cycle of stabilizer measurements Andersen2020 .
@@
    Fully Functional Quantum Computers (FFQC). Existing quantum algorithms and circuits consider the use of ancillary qubits, which play a similar role as cache memory in classical processors. However, for Fully Functional Quantum Computers, there will also be necessary the existence of Quantum Random Access Memories (qRAM) QuantumRam , and even long-term (error-corrected) quantum memories Terhal2015 .
We live currently in the NISQ era, while the community is still investing to keep advancing quantum computers to the next stages. To meet the PQEC era, we need to further explore error correction codes. It is important to consider that the number of physical qubits required by the error corrections codes to correct any error generally depends on how below the threshold the processor and operations are Andersen2020 ; suppressingQuantumErrors . The generation of quantum processors developed in this era could be considered a transition age towards a fault-tolerant quantum computer in which, in principle, there are no limitations in the size of the quantum circuits that can be reliably executed. It is important to remark that these codes rely on sufficiently trustworthy qubits and operations, not only quantum gates, but also very accurate measurements.
The challenges that we are facing today within the NISQ era are different from the ones we will be facing in the upcoming eras, and hence the benchmarks should reflect this. Given the wide diversity of technological platforms and paradigms for implementing quantum computers, it is likely that benchmarks in the NISQ era will focus on comparing the performance among different quantum computers within subsets of the broad existing variety. As quantum computing technology advances, it is likely that certain technological platforms will become dominant, particularly those that demonstrate scalability, fault tolerance, or economic viability for large-scale implementation. However, this convergence does not avoid the possibility of continued diversification. As seen in classical computing with the emergence of application-specific integrated circuits (ASICs), technological maturity can also drive specialization tailored to distinct tasks. Similarly, in quantum computing, different platforms may evolve to address specific application domains. In Sec. IV, we will discuss further what recommendations we suggest to properly face the different quantum technological development eras.
@@
  • Even though an $n$-qubit quantum state can in principle contain an exponential quantity of information in bits 10.1145/581771.581773 , the amount of classical information that can be retrieved by a projective measurement is limited to, at most, $n$ bits due to Holevo’s bound Maslov_2021 . As a consequence, if the algorithm requires to retrieve the full quantum state, i.e. to perform full tomography of the state, then an exponential number of measurements will be required.
@@
  • A quantum computer execution typically sticks to the Input-Process-Output (IPO) model, a rather general framework used in software engineering to describe the structure of an information processing program. Generally, a quantum computer needs to complete three different stages: the input or loading of information, the processing thereof, and the output or retrieval of relevant information. Most quantum algorithms adhere to this model PPP , where the input is state preparation or matrix exponentiation, and the output is just the quantum measurement (see Fig.1), adding extra steps to the classical algoritmic execution.
@@
III.2 A common language to discuss
The lack of a common and accepted standard for defining techniques and processes in quantum computing has led to the misuse of terminology. Even though there has been past attempts to classify the different types of benchmarks, and review approaches in the field, the terms quantum benchmarking, validation, verification, and testing have been loosely and inconsistently used in the literature. While we appreciate the groundwork laid by amico2023defining ; proctor2024benchmarkingquantumcomputers in establishing a foundation for benchmarking standards and properties, our approach builds upon not only insights from quantum computing research but also draws from the well-established and mature field of classical computing benchmarking. In addition, we support our proposal with numerous carefully analyzed examples, which help to illustrate and validate the concepts and classifications we introduce. In this subsection, we will propose a unified terminology for the main concepts used in the quantum benchmarking field.

### v4

@@ 2.2 What types of benchmarks are there?
Benchmarks can be classified according to multiple criteria, ranging from the objective of what portion of the behavior space they want to assess to the way the workload is defined. All of these categories fall under one of the following three different strategies when trying to assess the performance [7, 9]:
@@
  • Fixed-work workloads: This benchmarking strategy refers to the types of benchmarks that focus on measuring the amount of time required to perform a fixed amount of computational work, as seen in the case of the old classic Dhrystone and Whetstone benchmarks. These benchmarks compiled a list of programs to be executed, the former favouring floating point operations, while the latter focuses on integer performance analysis [14, 15].
@@
  • Fixed-time workloads: This benchmarking strategy refers to the types of benchmarks that focus on measuring the amount of computation performed within a fixed period, as seen in the case of the SLALOM benchmark. This benchmark runs an algorithm to calculate the radiosity, with the performance metric being the accuracy of the answer computed within a fixed time [44].
@@
  • Variable-work and variable-time workloads: This benchmarking strategy refers to the types of benchmarks that allow both the amount of computation performed and the time to vary. A great example is the HINT benchmark that aimed to find rational upper and lower bounds for $\int_{0}^{1}\frac{1-x}{1+x}.\,dx$. The assessment was later performed by calculating the quality improvements per second (QUIPS), i.e. dividing the quality of the solution by the execution time [38].
@@
In general, the performance of a system depends on the performance of multiple components of the system that are utilized during a certain operation. The problem with using fixed-work benchmarks is that they introduce an intrinsic performance bottleneck when measuring the increment of improving a single component of the system. Benchmarking models have been developed changing the workload type precisely to address this concern, and further discussions on overcoming issues with the fixed-work via using fixed-time benchmarks can be found in Ref. [45, 46]. However, benchmarks that allow for variability in both time and workload provide greater flexibility to evaluate the performance of different systems, determining which machine produces the highest-quality result in the shortest time possible.
@@
One approach to benchmarking classification is obtained by focusing on the level of implementation required from users[5, 9]:
@@
The specification-based benchmarks are sometimes called “pen and paper” benchmarks, and they have proved successful when there are considerable differences in the architectures on which the benchmarks will be implemented [47]. Still, they come at the cost of ad-hoc and expensive implementations from their users, as well as challenges with verifiability. On the other hand, kit-based benchmarks reduce the cost and time to run them by offering execution-ready implementation, which favoured them historically. Indeed, the stone aged benchmarks for classical computing are good examples of this, such as the Whestone, Dhrystone or the Linpack benchmarks [14, 15, 48].
Another approach to benchmarking classification is obtained by focusing on the workload [9, 7]:
@@
2.3 What makes a good metric
@@
From Lilja and Kounev [7, 9], we know that these can be typically classified into three groups: Counting how often a certain event occurs (e.g. number of operations processed in a given observation interval), duration of an interval (e.g. time taken to complete a single operation), and size of a parameter (e.g. precision obtained to compute a value). Some examples of basis performance metrics are the response time (i.e., time taken by a system to react to a request providing a respective response), throughput (i.e., rate at which requests are processed by a system), and utilization (i.e., fraction of time in which a resource is used).
Similarly to the quality attributes defined for a benchmark, there are properties proposed in the literature for an ideal performance metric [9, 7, 5]:
@@
We can also classify metrics depending on what they measure. Metrics that measure what was done, useful or not, have been called means-based metrics whereas end-based metrics measure what is actually accomplished. As an example, for a given program we could measure the floating point operations per second (FLOPS), or the total execution time. The former is a mean-based metric, while the latter is an end-based metric. Depending on the context, it could be argued that one is better than the other [7].
@@
Other well-known examples of classical metrics are clock rate, or million instructions per second (MIPS). The clock rate is repeatable (given a fixed processor), practical, consistent, and independent metric. However, given its limited scope, the clock rate is both non-linear and unreliable as a performance metric (a processor with a faster clock rate does not necessarily imply better overall performance. The clock rate ignores many important performance-relevant aspects such as how much computation is actually performed in each clock cycle as well as the interaction with the memory and I/O subsystems). On the other hand, MIPS is neither reliable, linear, or consistent (for the latter, take the example of different Instruction Set Architectures, such as RISC vs CISC [49]). Ultimately, we mentioned before FLOPS. Unfortunately, this metric is neither reliable, linear, consistent, nor independent (processor manufactures may use different rules for what is counted as floating-point operations, which makes the metric fail for consistency and independence).
@@
Finally, it is important to highlight two key points. First, the properties discussed in this work are applicable not only to full-stack benchmarks and metrics but also to individual components of the computing stack. Second, the seemingly straightforward task of measuring a property using a metric might be tough. The effect of the Simpsons’ paradox serves as an adage to the designers of benchmarks and metrics, calling out the possibility of reversing or disappearing the trend on the data when combining it [10].
2.4 Examples of classical benchmarks
Classical computing benchmarking has evolved to address multiple and distinct application domains, from machine learning to high-performance computing (HPC), each with domain-specific requirements. Experience with classical benchmarks has shown that defining benchmarks that remain relevant, fair, and consistent is a nontrivial challenge. For instance, the extensive use of LINPACK and its successor HPLinpack for ranking supercomputers has revealed that such benchmarks may fail to accurately reflect performance on real-world HPC applications, while also encouraging optimization strategies tailored to the benchmark rather than to relevant workloads [50, 51, 52, 53, 54]. Similarly, synthetic benchmarks such as Whetstone [14] have proven controversial. Although designed to measure floating-point performance, they can be misleading, as modern architectures may exploit cache effects and pipeline optimizations that are unrepresentative of real workloads. Historically, this issue was further compounded by the lack of consistent definitions of what constituted a “floating-point operation” across architectures and instruction sets, making cross-platform comparisons ambiguous in early benchmarking efforts. More generally, synthetic benchmarks often correlate poorly with practical application performance, limiting their relevance. These limitations highlight the need for continuous adaptation of benchmarking methodologies as technology evolves. This approach has been adopted by the Standard Performance Evaluation Corporation (SPEC), which regularly updates its benchmark suites and develops application-specific benchmarks targeting different workload classes, such as SPECviewperf for graphics workloads, SPECaccel for accelerator-based applications, and SPEC CPU for compute-intensive workloads [55, 56, 57]. These lessons from classical benchmarking motivate a careful, principle-based approach to quantum benchmarking, especially given the rapid evolution and heterogeneity of quantum hardware platforms.
Key points of the section: What can we learn from Classical benchmarking? What benchmarking is: In computing, a benchmark is a test (or set of tests) designed to compare the performance of one system against others. What performance means: Performance is defined as the useful work done within a specific time or using specific resources. Properties of a “Good” Benchmark: We adopt the SPEC definition, which outlines five essential qualities: Relevance: Correlates with user interests (speed, energy, accuracy). Reproducibility: Produces consistent results under the same configuration. Fairness: Allows competition without bias (e.g., using “Base” metrics for standard compilation rules vs. “Peak” metrics for optimized flexibility). Verifiability: Results can be independently confirmed; ideally includes self-validation. Usability: Easy and affordable to run. Types of Benchmarks: Benchmarks are classified by three main criteria: By Strategy (Time vs. Work): Fixed-work, fixed-time, and variable-work/variable-time. By Implementation Level: Specification-based, kit-based, hybrid. By Workload type: Synthetic, microbenchmarks, kernels and applications. Metric: A metric characterizes a performance property (count, duration, or size), and can be classified in two categories: means-based that measures what was done (e.g., FLOPS), and end-based that measures what was accomplished (e.g., total execution time). Following SPEC, a “good” metric should be: Practical: Efficient/polynomial time to compute. Repeatable: Deterministic results. Reliable: Consistently ranks systems (System A >> System B). Linear: Twice the value equals twice the performance. Consistent: Same units and definitions across systems. Lessons from History Adaptation is Key: Benchmarks must evolve. LINPACK is used for ranking supercomputers but often fails to reflect real-world HPC applications, encouraging optimization for the benchmark rather than the workload. Synthetic Risks: Benchmarks like Whetstone can be misleading on modern hardware because they do not account for modern cache effects or pipeline optimizations. No Silver Bullet: No single benchmark can evaluate all aspects of a computer. A comprehensive assessment requires a suite of benchmarks (like SPEC CPU or SPECviewperf) to model the full behavior space.
3 Quantum benchmarking
@@
Quantum computers are devices that leverage the principles of quantum mechanics to address challenges beyond the computational capacity of classical computers. Several physical implementations for quantum computers have been explored, each employing different technologies as qubits, such as trapped ion [58, 59], superconducting circuits [60, 61, 62], photonic [63, 64], and cold atoms [65, 66]. Each of these architectures has its own set of advantages and disadvantages, and research is ongoing to determine which approach is most suitable for scaling up quantum computers to a practical level. Additionally, various quantum computing paradigms have emerged, including annealing-based quantum computing, which relies on the adiabatic theorem to solve computational problems [67]; analog quantum computing, in which a controllable quantum device used to simulate other quantum systems [68]; digital quantum computing, which relies on the quantum circuit model for implementing general quantum algorithms [69]; and measurement-based quantum computing, where computation is performed through a sequence of measurements on a highly entangled state [70]. These four paradigms offer different perspectives on the practical applications of quantum computing.
@@
The diversity of technological platforms and paradigms used in the development of quantum computers introduces significant complexity in comparing the performance among different devices. For this reason, the search for a common benchmark that allows comparison of quantum computers performance across different technological platforms and paradigms poses a challenge to the scientific community. Despite these challenges, efforts have been made to develop benchmarks enabling comparison among diverse quantum computers. Primarily, these efforts focus on three types of relevant comparisons: comparing the performance of different quantum processors within the same computational paradigm (e.g., comparison of digital quantum computing platform providers such as Forest, Qiskit, ProjectQ [71], or comparisons between quantum processors [72, 73, 74, 75]); comparing different computational paradigms (e.g., between processors for digital quantum computing and quantum annealing or digital-analog quantum computing [76]); and finally, comparing quantum computers to classical computers (e.g. comparing digital processing against quantum annealing [77], empirical comparisons relying upon simplified Machine Learning (ML) tasks [78, 79] or even the claims on quantum supremacy [60, 80]). It is worth noting that these three comparisons are crucial for advancing scalable and reliable quantum technologies, but each of them demands a thorough analysis. Furthermore, a similar analysis could be expanded to include algorithmic comparisons. Looking ahead, benchmarking should be extended to cover other potential new aspects of a quantum computer that we have not even envisioned yet (e.g. qRAM [81] or other future components that we have not even thought about yet). During this work, we will restrict ourselves to the evaluation of the performance of quantum processors.
@@
Despite technological advances, we are still undergoing the development of the quantum computing foundations. As Blume-Kohout and Young stated [22], we are still more in the early stages of the quantum computing technology world rather than close to any exponential growth, which translates into challenges as we do not have clearly defined applications for these devices. What is more, we lack knowledge of how each the performance of each component will propagate and how the different quantum technology architectures will evolve. The current state is closely related to classical devices in the pre-1940 stages with the introduction of devices that were able to mechanize sequential series of calculations and procedures in binary code via the Bell Labs Model I, which served as an example of successfully applying Boolean logic to the design of computers [82]. Back then, computers were starting to be explored, in contrast to upcoming years such as 1965 when Moore published the famous paper that led to the establishment of the Moore’s law [83], and exponential growth took off. This contrasts with the “benchmarketing” approach that we sometimes see in the community when promoting the results of quantum computing experiments.
@@
3.1 Nuances of benchmarking quantum computers
There is lack of a definite computational paradigm, standard, platform, or even architecture for quantum computers. This is an eventual consequence of our limited knowledge so far about the real scope of quantum computing devices and platforms. Note that, in classical computing, regardless of the type of architecture (i.e. Princeton/Von Neumann vs. Harvard), there is a similar unified approach that allows for performance evaluation using well-established metrics, workloads, and even benchmarks. Indeed, it can be argued that differences in the architectures have been historically deceptive [84], and that devices can be easily compared assuming a similar common ground. Interestingly, this was not always the case. As mentioned before, back in the times of the Bell Labs Model I, every computer device was custom and hard to compare against each other. Benchmarking was not so extended and generalised as a recommended best practice until the 70s-80s.
@@ 3.1 Nuances of benchmarking quantum computers
    Noisy Intermediate-Scale Quantum (NISQ) [85], characterized by middle-size quantum processors with tens to hundreds of noisy physical qubits, substantially limiting the size of the quantum circuits that can be reliably executed. In the current state of NISQ era, it is controversial whether quantum processors have already reached the necessary technological development to achieve quantum advantage.
@@
    Partially Quantum Error Corrected (PQEC), when quantum processors errors are below a certain threshold that allows the application of error correction codes, such as the surface code [86]. When an insufficient number of physical qubits are available, we could employ a smaller code and partially correct errors up to a certain distance-code and number of errors per per cycle of stabilizer measurements [87].
@@
    Fully Functional Quantum Computers (FFQC). Existing quantum algorithms and circuits consider the use of ancillary qubits, which play a similar role as cache memory in classical processors. However, for Fully Functional Quantum Computers, there will also be necessary the existence of Quantum Random Access Memories (qRAM) [81], and even long-term (error-corrected) quantum memories [88].
We live currently in the NISQ era, while the community is still investing to keep advancing quantum computers to the next stages. To meet the PQEC era, we need to further explore error correction codes. It is important to consider that the number of physical qubits required by the error corrections codes to correct any error generally depends on how below the threshold the processor and operations are [87, 89]. The generation of quantum processors developed in this era could be considered a transition age towards a fault-tolerant quantum computer in which, in principle, there are no limitations in the size of the quantum circuits that can be reliably executed. It is important to remark that these codes rely on sufficiently trustworthy qubits and operations, not only quantum gates, but also very accurate measurements.
The challenges that we are facing today within the NISQ era are different from the ones we will be facing in the upcoming eras, and hence the benchmarks should reflect this. Given the wide diversity of technological platforms and paradigms for implementing quantum computers, it is likely that benchmarks in the NISQ era will focus on comparing the performance among different quantum computers within subsets of the broad existing variety. As quantum computing technology advances, it is likely that certain technological platforms will become dominant, particularly those that demonstrate scalability, fault tolerance, or economic viability for large-scale implementation. However, this convergence does not avoid the possibility of continued diversification. As seen in classical computing with the emergence of application-specific integrated circuits (ASICs), technological maturity can also drive specialization tailored to distinct tasks. Similarly, in quantum computing, different platforms may evolve to address specific application domains. In Sec. 4, we will discuss further what recommendations we suggest to properly face the different quantum technological development eras.
@@
  • Even though an $n$-qubit quantum state can in principle contain an exponential quantity of information in bits [90], the amount of classical information that can be retrieved by a projective measurement is limited to, at most, $n$ bits due to Holevo’s bound [91]. As a consequence, if the algorithm requires to retrieve the full quantum state, i.e. to perform full tomography of the state, then an exponential number of measurements will be required.
@@
  • A quantum computer execution typically sticks to the Input-Process-Output (IPO) model, a rather general framework used in software engineering to describe the structure of an information processing program. Generally, a quantum computer needs to complete three different stages: the input or loading of information, the processing thereof, and the output or retrieval of relevant information. Most quantum algorithms adhere to this model [92], where the input is state preparation or matrix exponentiation, and the output is just the quantum measurement (see Fig.1), adding extra steps to the classical algoritmic execution.
@@
3.2 Why platform heterogeneity precludes rigid numerical standards
In classical computing, heterogeneity among devices is mainly architectural, while the underlying operational model and logical abstractions remain consistent across platforms. In contrast, quantum computing is currently in a stage characterized by a high degree of platform diversity. Quantum platforms differ in several aspects, including the underlying physical technologies used to implement qubits, the set of native quantum operations they support, and the topology of the quantum hardware. As a consequence, performance-related properties such as gate execution times, qubit connectivity, and dominant error mechanisms exhibit differences across platforms, challenging the use of numerical thresholds to define performance acceptability. For this reason, specifying quantitative cutoffs for performance properties is inappropriate at this stage.
For instance, gate execution times span several orders of magnitude across platforms, ranging from microseconds to milliseconds, depending on the underlying physical technology. As a result, a metric that relies on the execution of deep quantum circuits may be feasible on platforms with fast gate operations, while becoming increasingly costly on platforms with significantly slower gates due to increased experimental runtime and accumulated decoherence. Consequently, defining what constitutes an acceptable or good execution time through a fixed numerical threshold does not carry the same meaning across different technologies. Likewise, quantum platforms differ in the native gate sets they support, ranging from architectures limited to one- and two-qubit gates to platforms enabling native multi-qubit operations. As a consequence, metrics based on gate counts or circuit depth may correspond to different computational effort across platforms, since a single native multi-qubit operation can replace an entire sequence of two-qubit gates in other architectures. Therefore, defining acceptable or meaningful thresholds on gate counts or circuit depth through fixed numerical values does not reflect equivalent operational cost across platforms. Similarly, qubit connectivity differs widely across platforms, ranging from nearest-neighbour connectivity to all-to-all connectivity. As a result, metrics that rely on interactions between distant qubits may incur different overheads depending on the connectivity of the quantum processor. Furthermore, error mechanisms vary qualitatively across technologies, with different platforms being dominated by distinct noise sources such as decoherence, control errors, crosstalk, or measurement noise. Therefore, metrics or benchmarks sensitive to particular error sources may reflect different performance constraints across platforms. As a result, fixed numerical thresholds on such metrics do not carry the same meaning across technologies.
These examples are only illustrative. In general, the heterogeneity of quantum platforms reinforces the difficulty of defining universal numerical thresholds for performance acceptability. Taken together, these considerations illustrate that the difficulty in defining rigid numerical standards for quantum benchmarking comes from the heterogeneity of quantum platforms. At the present stage of quantum hardware development, a guideline-based framework provides a more robust approach, allowing benchmarks and metrics to be evaluated in context without imposing premature numerical standardization.
3.3 A common language to discuss
The lack of a common and accepted standard for defining techniques and processes in quantum computing has led to the misuse of terminology. Even though there has been past attempts to classify the different types of benchmarks, and review approaches in the field [93, 94], the terms quantum benchmarking, validation, verification, and testing have been loosely and inconsistently used in the literature. While we appreciate the groundwork laid by [25, 95] in establishing a foundation for benchmarking standards and properties, our approach builds upon not only insights from quantum computing research but also draws from the well-established and mature field of classical computing benchmarking. In addition, we support our proposal with numerous carefully analyzed examples, which help to illustrate and validate the concepts and classifications we introduce. In this subsection, we will propose a unified terminology for the main concepts used in the quantum benchmarking field.